# Polynomial time auditing of statistical subgroup fairness for Gaussian data

### 27 Jan 2024 v1

@@ 3.2 The Hardness of Auditing Under Gaussian Data
3.2.1 Approximation in Multiplicative Form
Suppose an auditing algorithm is guaranteed to return us a $\gamma^{\prime}$-unfair certificate (a halfspace) give a $\gamma$-unfair classifier $c$, where $\gamma^{\prime}\leq\gamma\leq 1$, we can show that $\gamma^{\prime}$ can never be close to $\gamma$. We can show that $\gamma^{\prime}=\Omega(\gamma/\mathrm{poly}(d))$ only if cLWE can be solved in $O(\mathrm{poly}(d))$ time.
@@ 3.2.1 Approximation in Multiplicative Form
Under Assumption 2.6, for any $d\in\mathbb{N}$, any constants $\alpha\in(0,1),\beta\in\mathbb{R}_{+}$, and any $\log^{\beta}d\leq k\leq cd$ where $c$ is a sufficiently small constant, there is not algorithm that runs in time $d^{O(k^{\alpha})}$ and distinguishes between the following two cases of a joint distribution $\mathcal{D}$ of $(\mathbf{x},c(\mathbf{x}))$ supported on $\mathbb{R}^{d}\times\{-1,+1\}$ with marginal $\mathcal{D}_{\mathbf{x}}=\mathcal{N}(0,\mathrm{I})$, with $d^{-O(k^{\alpha})}$ advantage:
@@
    Alternative Hypothesis: There exists non-negligible unfair halfspace subgroups, i.e., $\exists h\in\mathcal{H}^{d},\Pr_{\mathcal{D}}\{h(\mathbf{x})=1\}|d_{\mathcal{D}}(c,h)|=\Omega(1/\sqrt{k\log d})$.
@@
    Null Hypothesis: All halfspace subgroups are absolutely fair, i.e., $\Pr_{\mathcal{D}}\{h(\mathbf{x})=1\}|d_{\mathcal{D}}(c,h)|=0,\forall h\in\mathcal{H}^{d}$.
@@
$=$ $\underbrace{\Pr\{\neg c\}(\Pr\{\neg h_{1}\}+\Pr\{\neg h_{2}\})+\Pr\{c\}(\Pr\{h_{1}\}+\Pr\{h_{2}\})}_{I_{1}}-(\underbrace{\Pr\{c(\mathbf{x})\neq h_{1}(\mathbf{x})\}+\Pr\{c(\mathbf{x})\neq h_{2}(\mathbf{x})\}}_{I_{2}})$
@@
where (i) is because $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=-1\},\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=-1\},\{\mathbf{x}\in B\}$ are pairwise disjoint and their union equals to $\mathbb{R}^{d}$, $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=1\},\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=1\ \cap\ \mathbf{x}\notin B\}$ are disjoint and their union equals to $\mathbb{R}^{d}$, also, $\{\mathbf{x}\in B\}\subset\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=1\}$ by definition.
@@
$I_{2}=$ $\Pr\{c(\mathbf{x})\neq h_{1}(\mathbf{x})\ \cap\ \mathbf{x}\notin B\}+\Pr\{c(\mathbf{x})\neq h_{2}(\mathbf{x})\ \cap\ \mathbf{x}\notin B\}+2\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}$
$=$ $\Pr\{\mathbf{x}\notin B\}+2\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}$
$=$ $1+\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}-\Pr\{c(\mathbf{x})=1\ \cap\ \mathbf{x}\in B\}$
$=$ $1-\Pr\{\mathbf{x}\in B\}(1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\}$
@@
$I_{1}-I_{2}=$ $\Pr\{\mathbf{x}\in B\}(1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\})$
$=$ $\Omega(T)(1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\})$ (12)
where the last equation is because $\mathbf{s}^{\top}\mathbf{x}\sim\mathcal{N}(0,1)$, hence, $\Pr\{\mathbf{x}\in B\}=\Pr\{\mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]\}=\Omega(T)$. Thus, it suffices to show $1-2\Pr\{c(\mathbf{x})=-1\ |\ \mathbf{x}\in B\}=\Omega(1)$.
Again, by the definition of the Alternative case distribution from the LWE problem, we know that $c(\mathbf{x})=-1$ only if $\mod_{T}(\mathbf{s}^{\top}\mathbf{x}+z)>T/2$, which in turn happens only if $\mathbf{s}^{\top}\mathbf{x}+z>T/2$ or $\mathbf{s}^{\top}\mathbf{x}+z<0$. For $\mathbf{x}\in B$, we have $\mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]$, therefore $c(\mathbf{x})=-1$ only if $\left|z\right|\geq T/6$. Notice that $z\sim\mathcal{N}(0,\sigma^{2})$ and Proposition 3.8 states that the LWE problem is hard for any fixed constant $\kappa\in\mathbb{N}$ and $\sigma\geq k^{-\kappa}$. Given the constant $\beta\in\mathbb{R}_{+}$ in this theorem, we can take $\kappa=\left\lceil 1/2\beta+1/2+1\right\rceil$, which is a fixed constant. Then, by Proposition 3.8, the LWE problem is hard for $\sigma=k^{-\kappa}\leq 1(k^{3/2}\sqrt{\log d})=o(T)$. Therefore, by Gaussian tail bound, we have
@@
Plugging the above back to equation (12), we can conclude that
@@
Thus, either $h=h_{1}$ or $h=h_{2}$ must satisfies $\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|=\Omega(T)$, which completes the proof for the alternative hypothesis case.
@@
It immediately indicates the following corollary.
Corollary 3.10.
Given Assumption 2.6, for any constants $\alpha\in(0,1),\log^{\beta}d\leq k\leq cd$ where $c$ is a sufficiently small constant, there is no polynomial-time $1/\mathrm{poly}(d)$-approximation algorithms for auditing halfspace subgroups under Gaussian marginal.
@@
For the alternative hypothesis case as described in Theorem 3.9, given a $1/\sqrt{k\log d}$-unfair collection of halfspace subgroups, we run such an algorithm to obtain a $\delta/\sqrt{k\log d}$-unfair certificate, i.e., a halfspace $h$ such that $\Pr_{\mathbf{x}\sim\mathcal{N}}\{h(\mathbf{x})=1\}|d_{\mathcal{N}}(c,h)|=\Omega(\delta/\sqrt{k\log d})$. Notice that by Hoeffding Bound, with $d^{O(k^{\alpha})}$ examples from the distribution constructed in the alternative hypothesis case, we can verify that the empirical estimation of $\Pr_{\mathbf{x}\sim\mathcal{N}}\{h(\mathbf{x})=1\}|d_{\mathcal{N}}(c,h)|$ is $d^{-O(k^{\alpha})}$-close to $\Omega(\delta/\sqrt{k\log d})$ with high probability.
@@
For the null hypothesis case, with the same argument, we can verify there is no $d^{-O(k^{\alpha})}$-unfair subgroup with high probability given $d^{O(k^{\alpha})}$ examples from the distribution in the null hypothesis case.
Therefore, the algorithm can distinguish the alternative hypothesis case from the null hypothesis case in Theorem 3.9 unless $\delta=d^{-O(k^{\alpha})}$, which implies the desired result. ∎
3.2.2 Approximation in Additive Form
According to Theorem 3.1, how well we can approximate the most unfair subgroup is controlled by the parameter $\epsilon$. Unfortunately, it turns out, even under a Gaussian distribution, $\epsilon$ cannot be as small as $1/\mathrm{poly}(d)$.
A recent result [17] showed that learning halfspaces with an additive error $O(1/\sqrt{\log d})$ is as least as hard as the problem of continuous Learning With Error (cLWE) under Gaussian distribution.
In particular, they showed an efficient way to generate $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{-1,+1\}$ from the data, $(\mathbf{x},y^{\prime})\in\mathbb{R}^{d}\times\mathbb{R}_{T}$, sampled from the cLWE structured Gaussian distribution such that the properties of both alternative and null hypothesis are preserved. In particular, under the distribution from the alternative hypotheses of cLWE, they claim that at least one of the two halfspaces, $\mathbf{s}^{\top}\mathbf{x}\geq T/6$ and $\mathbf{s}^{\top}\mathbf{x}\leq T/3$, must have classification error less than $1/2-\Omega(T)$. Hence, if one can learn halfspaces, they can also distinguish the two cases in cLWE.
We can reduce cLWE to auditing by generating data in the same way as [17]. For the null hypothesis, we have that $\Pr_{\mathbf{x}\in\mathcal{N}}\{h\}d_{\mathcal{N}}(c,h)=0,\forall h\in\mathcal{H}^{d}$ because $c(\mathbf{x})$ is independent from each $h\in\mathcal{H}^{d}$. For the alternative hypothesis, the same two halfspaces, $\mathbf{s}^{\top}\mathbf{x}\geq T/6$ and $\mathbf{s}^{\top}\mathbf{x}\leq T/3$, will result to a unfairness level distinguishable from 00 based on Lemma 3.6. Therefore, being able to audit $h_{1}$ and $-h_{2}$ such that $h_{1}\in\mathcal{H}_{T/6}^{\mathcal{N}}$ and $h_{2}\in\mathcal{H}_{-T/3}^{\mathcal{N}}$ to an $\epsilon$ error will allow us to solve cLWE problem. This observation yields the following corollary.
Corollary 3.11.
@@ 3.2.2 Approximation in Additive Form
Given Assumption 2.6, for any constants $\alpha\in(0,2),\beta>1/2$ and any $C/\sqrt{d\log d}\leq\epsilon\leq 1/\log^{\beta}d$ where $C$ is a sufficiently large constant, there is no algorithm running in time $d^{O(\epsilon\sqrt{\log d})^{\alpha}}$ that can recover a subgroup represented by a halfspace on $\mathbb{R}^{d}$ with Gaussian marginals with unfairness within an additive $\epsilon$ of maximum.
For details about the reduction, we refer to the proof of [17].
@@
To the best of our knowledge, there does not exist any PTAS approach for properly learning general halfspaces in agnostic model with guarantees of additive error close to $O(1/\sqrt{\log d})$. However, in the next section, we will show that if we restrict out attention to just homogeneous halfspaces under a standard normal distribution, it is possible to achieve additive error of $O(1/\log^{1/C}d)$ for some constant $C>2$.
@@ 17      $\mu\leftarrow\mu+(b-a)/n$;
However, if we restrict our auditing subjects to the class of homogeneous halfspaces, [18] proposed an agnostic learning PTAS for homogeneous halfspaces under Gaussian data. That is, we only audit for subgroups with probability mass $1/2$.
@@
Let $\mathcal{D}$ be a distribution on labeled examples $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{-1,+1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(0,\mathrm{I})$. There exists an algorithm that, given $\epsilon,\delta>0$, and $N=d^{O(1/\epsilon^{4})}\mathrm{poly}(1/\epsilon)\log(1/\delta)$ i.i.d. samples from $\mathcal{D}$, the algorithm runs in time $\mathrm{poly}(N)+(1/\epsilon)^{O(1/\epsilon^{6})}\log(1/\delta)$, and computes a homogeneous halfspace $h_{\mathbf{v}}$ such that, with probability at least $1-\delta$, it holds $\Pr_{\mathcal{D}}\{y\neq h_{\mathbf{v}}(\mathbf{x})\}\leq\mathrm{opt}+\epsilon$.
@@
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{N}(0,\mathrm{I})$ and a collection of halfspaces $\mathcal{H}_{1/2}^{\mathcal{N}}$ over $\mathbb{R}^{d}$, there exists an auditing algorithm that takes $\delta>0$ and $N=d^{O(1/\epsilon^{4})}\mathrm{poly}(1/\epsilon)\log(1/\delta)$ labeled i.i.d. example from $\mathcal{N}(0,\mathrm{I})$ in the form of $(\mathbf{x},c(\mathbf{x}))$, runs in time $\mathrm{poly}(N)+(1/\epsilon)^{O(1/\epsilon^{6})}\log(1/\delta)$, and returns a halfspace $h^{\prime}$ as a certificate such that $\Pr_{\mathbf{x}\sim\mathcal{D}}\{h^{\prime}\}=1/2$ and
$\left|d_{\mathcal{N}}(c,h^{\prime})\right|\geq\max_{h\in\mathcal{H}_{1/2}^{\mathcal{N}}}\left|d_{\mathcal{N}}(c,h)\right|-\epsilon$
@@
Simply running algorithm 1 with the same set of parameters except that $\mathcal{D}=\mathcal{N}(0,\mathrm{I})$, $n=1$, $a=b=1/2$ and the oracle being as described by lemma 4.2 will give us the desired results. ∎

### 29 Feb 2024 v2

@@ 4 The Hardness of Auditing Under Gaussian Data
We first show it is computationally hard to distinguish between halfspace subgroups that are evenly fair and halfspace subgroups among which there exists a slightly unfair subgroup.
@@
Under Assumption 2.7, for any $d\in\mathbb{N}$, any constants $\alpha\in(0,1),\beta\in\mathbb{R}_{+}$, and any $\log^{\beta}d\leq k\leq cd$ where $c$ is a sufficiently small constant, there is no algorithm that runs in time $d^{O(k^{\alpha})}$ and distinguishes between the following two cases of a joint distribution $\mathcal{D}$ of $(\mathbf{x},c(\mathbf{x}))$ supported on $\mathbb{R}^{d}\times\{-1,+1\}$ with marginal $\mathcal{D}_{\mathbf{x}}=\mathcal{N}(0,\mathrm{I})$, with $d^{-O(k^{\alpha})}$ advantage:
@@
    Alternative Hypothesis: There exist non-negligibly unfair halfspace subgroups, specifically $\exists h\in\mathcal{H}^{d},\Pr_{\mathcal{D}}\{h(\mathbf{x})=1\}|d_{\mathcal{D}}(c,h)|=\Omega(1/\sqrt{k\log d})$.
@@
    Null Hypothesis: All halfspace subgroups are perfectly fair, i.e., $\Pr_{\mathcal{D}}\{h(\mathbf{x})=1\}|d_{\mathcal{D}}(c,h)|=0,\forall h\in\mathcal{H}^{d}$.
@@
$=$ $\underbrace{\Pr\{\neg c\}(\Pr\{\neg h_{1}\}+\Pr\{\neg h_{2}\})+\Pr\{c\}(\Pr\{h_{1}\}+\Pr\{h_{2}\})}_{I_{1}}$
$-(\underbrace{\Pr\{c(\mathbf{x})=h_{1}(\mathbf{x})\}+\Pr\{c(\mathbf{x})=h_{2}(\mathbf{x})\}}_{I_{2}})$
@@
$=$ $1+\Pr\{\mathbf{x}\in B\}(2\Pr\{c(\mathbf{x})=1\}-1)$
where (i) is because $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=-1\},\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=-1\},\{\mathbf{x}\in B\}$ are pairwise disjoint and their union equals to $\mathbb{R}^{d}$, $\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{1}(\mathbf{x})=1\},\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=1\ \cap\ \mathbf{x}\notin B\}$ are disjoint and their union equals to $\mathbb{R}^{d}$; and since $\{\mathbf{x}\in B\}\subset\{\mathbf{x}\in\mathbb{R}^{d}\ |\ h_{2}(\mathbf{x})=1\}$ by definition, $\{\mathbf{x}\in B\}=\{\mathbf{x}\in B\ |\ h_{2}(\mathbf{x})=1\}$.
@@
$I_{2}=$ $\Pr\{c(\mathbf{x})=h_{1}(\mathbf{x})\ \cap\ \mathbf{x}\notin B\}+\Pr\{c(\mathbf{x})=h_{2}(\mathbf{x})\ \cap\ \mathbf{x}\notin B\}+2\Pr\{c(\mathbf{x})=1\ \cap\ \mathbf{x}\in B\}$
$=$ $\Pr\{\mathbf{x}\notin B\}+2\Pr\{c(\mathbf{x})=1\ \cap\ \mathbf{x}\in B\}$
$=$ $1+\Pr\{c(\mathbf{x})=1\ \cap\ \mathbf{x}\in B\}-\Pr\{c(\mathbf{x})=-1\ \cap\ \mathbf{x}\in B\}$
$=$ $1-\Pr\{\mathbf{x}\in B\}(1-2\Pr\{c(\mathbf{x})=1\ |\ \mathbf{x}\in B\}$
@@
$I_{1}-I_{2}=$ $2\Pr\{\mathbf{x}\in B\}(\Pr\{c(\mathbf{x})=1\}-\Pr\{c(\mathbf{x})=1\ |\ \mathbf{x}\in B\})$
$=$ $\Omega(T)(1/2-\Pr\{c(\mathbf{x})=1\ |\ \mathbf{x}\in B\})$ (12)
where the last equation is because $\mathbf{s}^{\top}\mathbf{x}\sim\mathcal{N}(0,1)$, hence, $\Pr\{\mathbf{x}\in B\}=\Pr\{\mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]\}=\Omega(T)$. Since we were only concerned with showing $|I_{1}-I_{2}|$ is large, it suffices to show $\Pr\{c(\mathbf{x})=1\ |\ \mathbf{x}\in B\}-1/2=\Omega(1)$.
For $\mathbf{x}\in B$, we have $\mathbf{s}^{\top}\mathbf{x}\in[T/6,T/3]$, therefore $c(\mathbf{x})=-1$ only if $\left|z\right|\geq T/6$. Notice that $z\sim\mathcal{N}(0,\sigma^{2})$ and Proposition 4.1 states that the LWE problem is hard for any fixed constant $\kappa\in\mathbb{N}$ and $\sigma\geq k^{-\kappa}$. Given the constant $\beta\in\mathbb{R}_{+}$ in this theorem, we can take $\kappa=\left\lceil 1/2\beta+1/2+1\right\rceil$, which is a fixed constant. Then, by Proposition 4.1, the LWE problem is hard for $\sigma=k^{-\kappa}\leq 1/(k^{3/2}\sqrt{\log d})=o(T)$. Therefore, by a Gaussian tail bound, we have
@@
Plugging the above back into equation (12), we can conclude that
@@
Thus, either $h=h_{1}$ or $h=h_{2}$ must satisfiy $\Pr_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\{h(\mathbf{x})=1\}\left|d_{\mathcal{D}}(c,h)\right|=\Omega(T)$, which completes the proof for the alternative hypothesis case.
@@
Suppose an auditing algorithm is guaranteed to return us a $\gamma^{\prime}$-unfair certificate (a halfspace) give a $\gamma$-unfair classifier $c$, where $\gamma^{\prime}\leq\gamma\leq 1$, the following corollaries show that $\gamma^{\prime}$ can never be close to $\gamma$.
Corollary 4.3 (multiplicative form).
Given Assumption 2.7, there is no polynomial-time $1/\mathrm{poly}(d)$-approximation algorithm for constructive auditing for halfspace subgroups under Gaussian marginals in $\mathbb{R}^{d}$.
@@
For the alternative hypothesis case as described in Theorem 4.2, given a $1/\sqrt{k\log d}$-unfair collection of halfspace subgroups, we run such an algorithm to obtain a $\delta/\sqrt{k\log d}$-unfair certificate, i.e., a halfspace $h$ such that $\Pr_{\mathbf{x}\sim\mathcal{N}}\{h(\mathbf{x})=1\}|d_{\mathcal{N}}(c,h)|\geq\delta/\sqrt{k\log d}$. By the Hoeffding Bound, we can verify that the empirical estimation of $\Pr_{\mathbf{x}\sim\mathcal{N}}\{h(\mathbf{x})=1\}|d_{\mathcal{N}}(c,h)|$ is $\varepsilon_{1}$-close to $\delta/\sqrt{k\log d}$ with high probability by drawing $O(1/\varepsilon_{1}^{2})$ examples from the distribution constructed in the alternative hypothesis case.
@@
For the null hypothesis case, with the same argument, we can verify there is no $\varepsilon_{2}$-unfair subgroup with high probability given $O(1/\varepsilon_{2}^{2})$ examples from the distribution in the null hypothesis case.
However, we are given that no algorithm can distinguish the two cases in Theorem 4.2 with $d^{-O(k^{\alpha})}$ advantage while running in time $d^{O(k^{\alpha})}$. Thus, we cannot have
$\frac{\delta}{\sqrt{k\log d}}-\varepsilon_{1}-\varepsilon_{2}=d^{-O(k^{\alpha})}$
for any $\varepsilon_{1},\varepsilon_{2}\leq d^{-O(k^{\alpha})}$ and any $\log^{\beta}d\leq k\leq cd$. Therefore, we can infer that $\delta$ cannot be as large as $d^{-O(k^{\alpha})}\sqrt{k\log d}=d^{-O(k^{\alpha})}$, which implies the desired result, since $k\geq\log^{\beta}d$ and hence $\delta<1/d^{\log^{\alpha\beta}d}$ which is negligible. ∎
Corollary 4.4 (additive form).
@@
Given Assumption 2.7, for any constants $\alpha\in(0,1),\beta\in\mathbb{R}_{+}$, and any $C/\sqrt{d\log d}\leq\epsilon\leq c^{\prime}/\log^{(1+\beta)/2}d$ where $C$ is a sufficiently large constant and $c^{\prime}$ is a sufficiently small constant, no auditing algorithm can return a unfair certificate for halfspace subgroups in $\mathbb{R}^{d}$ with an additive error $\epsilon$ under Gaussian marginals and runs in time $d^{O(1/(\epsilon^{2}\log d)^{\alpha})}$.
Proof.
Suppose there exists an auditing algorithm that guarantees to return a $\gamma-\epsilon$-unfair certificate given a $\gamma$-unfair collection of halfspace subgroup and access to data with Gaussian marginal, where $\epsilon\in(0,1)$.
Similar to the proof of Corollary 4.3, given a $1/\sqrt{k\log d}$-unfair collection of halfspace subgroups, we run such an algorithm to obtain a $(1/\sqrt{k\log d}-\epsilon)$-unfair certificate. Observe that, if $\epsilon=c^{\prime}/\sqrt{k\log d}$ for some sufficiently small constant $c^{\prime}$, we can solve the testing problem in Theorem 4.2 within time $d^{O(k^{\alpha})}$ by running this algorithm as well as drawing enough examples to estimate the unfairness of the returned certificates from the two cases respectively. On the other hand, given $\epsilon=c^{\prime}/\sqrt{k\log d}$, we can rewrite $d^{O(k^{\alpha})}=d^{O(1/(\epsilon^{2}\log d)^{\alpha})}$.
However, Theorem 4.2 tells that the above case is impossible for any $C/\sqrt{d\log d}\leq\epsilon\leq c^{\prime}/\log^{(1+\beta)/2}d$, where $C$ is a sufficiently large constant. ∎
Besides the general general auditing problem, we also consider the “non-constructive auditing” problem as in Definition 2.4, where the algorithm is only required to tell if there exists an unfair subgroup without returning the unfair certificate. Actually, it turns out any non-constructive auditing algorithm can distinguish the two cases in Theorem 4.2.
Corollary 4.5 (non-constructive auditing is hard).
Given Assumption 2.7, for any constants $\alpha\in(0,1),\beta\in\mathbb{R}_{+}$, and any $C/\sqrt{d\log d}\leq\epsilon\leq c^{\prime}/\log^{(1+\beta)/2}d$ where $C$ is a sufficiently large constant and $c^{\prime}$ is a sufficiently small constant, no auditing algorithm can tell if there exists a unfair certificate for halfspace subgroups in $\mathbb{R}^{d}$ with
  • an additive error $\epsilon$ under Gaussian marginals and running in time $d^{O(1/(\epsilon^{2}\log d)^{\alpha})}$.
  • or a multiplicative approximation factor of $1/\mathrm{poly}(d)$ and running in polynomial time.
Proof.
Suppose there exists an auditing algorithm that can either tell if a $\delta\gamma$-unfair certificate or a $\gamma-\epsilon$-unfair certificate exists given a $\gamma$-unfair collection of halfspace subgroup and access to data with Gaussian marginal, where $\delta,\epsilon\in(0,1)$. With the same argument as that of Corollary 4.3 and 4.4, we can achieve the desired results. ∎
@@
To the best of our knowledge, there does not exist any PTAS for properly learning general halfspaces in agnostic model with guarantees of additive error close to $O(1/\sqrt{\log d})$. However, in the next section, we will show that if we restrict out attention to just homogeneous halfspaces under a standard normal distribution, it is possible to achieve additive error of $O(1/\log^{1/C}d)$ for some constant $C>2$.
@@ 17      $\mu\leftarrow\mu+(b-a)/n$;
However, if we restrict our audit to the class of homogeneous halfspaces, Diakonikolas et al. [18] proposed an agnostic learning PTAS for homogeneous halfspaces under Gaussian data. That is, we only audit for subgroups with probability mass $1/2$.
@@
Let $\mathcal{D}$ be a distribution on labeled examples $(\mathbf{x},y)\in\mathbb{R}^{d}\times\{-1,+1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(0,\mathrm{I})$. There exists an algorithm that, given $\tau,\epsilon,\delta>0$, and $N=d^{\mathrm{poly}(1/\tau)}\mathrm{poly}(1/\epsilon)\log(1/\delta)$ i.i.d. samples from $\mathcal{D}$, the algorithm runs in time $\mathrm{poly}(N,d)$, and computes a halfspace $h_{\mathbf{v}}$ such that, with probability at least $1-\delta$, it holds $\Pr_{\mathcal{D}}\{y\neq h_{\mathbf{v}}(\mathbf{x})\}\leq(1+\tau)\min_{h\in\mathcal{H}_{1/2}^{\mathcal{N}}}\Pr_{\mathcal{D}}\{y\neq h(\mathbf{x})\}+\epsilon$.
@@
Given any binary classifier $c:\mathbb{R}^{d}\rightarrow\{-1,+1\}$, a data distribution $\mathcal{N}(0,\mathrm{I})$ and a collection of halfspaces $\mathcal{H}_{1/2}^{\mathcal{N}}$ over $\mathbb{R}^{d}$, there exists an auditing algorithm that takes $\epsilon,\delta>0$ and $N=d^{\mathrm{poly(1/\epsilon)}}\mathrm{poly}(1/\epsilon)\log(1/\delta)$ labeled i.i.d. example from $\mathcal{N}(0,\mathrm{I})$ in the form of $(\mathbf{x},c(\mathbf{x}))$, runs in time $\mathrm{poly}(N,d)$, and returns a halfspace $h^{\prime}$ as a certificate such that $\Pr_{\mathbf{x}\sim\mathcal{D}}\{h^{\prime}\}=1/2$ and
$\left|d_{\mathcal{N}}(c,h^{\prime})\right|\geq\max_{h\in\mathcal{H}_{1/2}^{\mathcal{N}}}\left|d_{\mathcal{N}}(c,h)\right|-2\epsilon$
@@
Simply running Algorithm 1 with the same set of parameters except that $\mathcal{D}=\mathcal{N}(0,\mathrm{I})$, $n=1$, $a=b=1/2$ and the oracle being as described by Lemma 5.2 for $\tau=\epsilon$ will give us the desired results. ∎